Find the solution to the system of equations
3x+2y=22
9x-8y=-4

Answer by oscargut(2103)

Find the solution to the system of equations
3x+2y=22
9x-8y=-4
-9x-6y=-66
9x-8y=-4

-14y = -70
Answer: y = 5, x = 4

If you have more questions i can see them if you contact me at: mthman@gmail.com
Thanks